OpenStudy (anonymous):

Verify that the functions f and g, are inverses of each other by showing that f(g(x))=x and g(f(x))=x. Graph both functions on the same graph. Please show all work. f(x)= 6/1-x g(x)= 1-6/x

OpenStudy (anonymous):

\[f(g(x))=\frac{6}{1-(1-\frac{6}{x})}\] \[g(f(x))=1-\frac{6}{\frac{6}{1-x}}\]

OpenStudy (anonymous):

can you graph it please

OpenStudy (anonymous):

Try out a few points: x | -5 | -2 | -1 | 0 | 1 | 2 | 3 | 4 | 7 f(x) | 1 | 2 | 3 | 6 | undefined | -6 | -3 | -2 | -1

OpenStudy (anonymous):

how did you get this answer?

OpenStudy (anonymous):

Hey vood- is there any way to get these on the same graph??